Find the nature of the planar section of the conicoid x²/3-y²/4=z by the plane
x+2y−z = 6
1
Expert's answer
2020-02-19T10:53:35-0500

"\frac{x^2}{3}-\frac{y^2}{4}=z\\x+2y-z=6"

We will find "z"

"z=x+2y-6"

"\frac{x^2}{3}-\frac{y^2}{4}=x+2y-6\\"

multiplying by 12

"4x^2-3y^2-12x-24y=-72\\
4(x^2-3x)-3(y^2+8y)=-72\\
4(x-\frac{3}{2})^2-3(y+4)^2=-72+4\cdot\frac{9}{4}-3\cdot16\\
4(x-\frac{3}{2})-3(y+4)^2=-111\\"

multiplying by "111"

"\frac{4(x-\frac{3}{2})}{111}-\frac{3(y+4)^2}{111}=-1\\
\frac{(x-\frac{3}{2})}{\frac{111}{4}}-\frac{(y+4)^2}{{\frac{111}{3}}}=-1\\"

hyperbola, conjugated hyperbola "\frac{(x-\frac{3}{2})}{\frac{111}{4}}-\frac{(y+4)^2}{{\frac{111}{3}}}=1\\" ,

 center "(\frac{3}{2};-4)" ,

semi-axes "a=\sqrt{\frac{111}{4}}, b=\sqrt{\frac{111}{3}}" .